# Teleportation=Translation: Continuous recovery of black hole information

### v2

@@ 4.1 Verification in Half-Sided Modular Inclusions
Our construction builds a strongly continuous path $\tilde{U}(s)$ that interpolates from the identity to this discrete shift, i.e., $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\Gamma}$. Assuming that this path is generated by a scale-independent operator $\tilde{G}$ (i.e., $\tilde{U}(s)\approx e^{-is\tilde{G}}$), the endpoint condition yields:
$e^{-i\tilde{G}}=\tilde{U}(1)=U_{\Gamma}=e^{-2iP}.$ (19)
This exact result in the HSMI limit strongly motivates the identification $\tilde{G}=2P$ for the general case. The factor of 2 in this relationship is dictated by the underlying modular structure. Physically, the canonical shift $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is composed of two consecutive modular conjugations. Since each modular conjugation effectively reverses the direction of the modular flow (analogous to a discrete time-reversal operation), combining two such operations results in a net translation that is double the elementary geometric shift. This justifies the factor of 2 in our operator identity.
4.2 Rigorous Derivation via Modular Perturbation Theory
Our goal is to prove the operator identity $\tilde{G}=2P$ in full generality (i.e., without relying on the special symmetry of the HSMI case). Since both $\tilde{G}$ and $2P$ are unbounded self-adjoint operators, a rigorous proof requires showing that they coincide as closed operators. We achieve this by employing the modular perturbation theory of Araki, Connes, and Kosaki (see, e.g., Araki:1976zv ; Connes80 ; Kosaki84 ; Kosaki84UC ), and then invoking Nelson’s theorem Nelson59 to establish essential self-adjointness.
@@ 4.2 Rigorous Derivation via Modular Perturbation Theory
For the canonical interpolation path $\tilde{\mathcal{N}}(s)$ (with $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}(1)=\tilde{\mathcal{N}}$) defined via the Connes–Takesaki cocycle derivative $D\phi_{1}:D\phi_{0}]_{t}$, the modular Hamiltonian $K(s)=K_{\tilde{\mathcal{N}}(s)}$ satisfies the following first-order perturbation relation at $s=0$:
@@
Let $\phi_{0}$ and $\phi_{1}$ be the faithful normal weights on $\tilde{\mathcal{M}}$ associated with the algebras $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$, respectively (such that $\Delta_{\tilde{\mathcal{M}}}=e^{-K_{\tilde{\mathcal{M}}}}$ and $\Delta_{\tilde{\mathcal{N}}}=e^{-K_{\tilde{\mathcal{N}}}}$). By Connes–Takesaki theory, the Radon–Nikodym cocycle relating these two weights is defined as:
$D\phi_{1}:D\phi_{0}]_{t}\;:=\;\Delta(\phi_{1}\,|\,\phi_{0})^{\,it}~,$ (21)
where $\Delta(\phi_{1}\,|\,\phi_{0})$ is the relative modular operator. The canonical interpolated weight $\phi_{s}$ (for $0<s<1$) is defined via the analytic scaling of this cocycle:
$D\phi_{s}:D\phi_{0}]_{t}\;=\;[D\phi_{1}:D\phi_{0}]_{\,st}~.$ (22)
This relation defines the path of weights and the corresponding algebras $\tilde{\mathcal{N}}(s)$ (via the range of the associated conditional expectations). Equivalently, the modular operator $\Delta_{s}=e^{-K(s)}$ for $\phi_{s}$ satisfies the relation:
$\Delta_{s}^{\,it}\;=\;\Delta_{0}^{\,it}\;\big[D\phi_{1}:D\phi_{0}\big]_{st}\,,\qquad\forall\,t\in\mathbb{R}~,$ (23)
where $\Delta_{0}=\Delta_{\tilde{\mathcal{M}}}$.
Differentiating both sides of (23) with respect to $s$ at $s=0$ (utilizing the differentiability guaranteed by the analytic construction) and considering the generator term, we observe:
$\frac{d}{ds}\Delta_{s}^{\,it}\Big|_{s=0}\;=\;\Delta_{0}^{\,it}\frac{d}{ds}\big[D\phi_{1}:D\phi_{0}\big]_{st}\Big|_{s=0}~.$ (24)
@@
Identifying the generator of the left-hand side of (24) similarly as $itK^{\prime}(0)$ (in the interaction picture defined by $\Delta_{0}$), we conclude:
$K^{\prime}(0)\;=\;K_{\tilde{N}}-K_{\tilde{M}}~,$ (26)
which is exactly the claimed result (noting $K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}}=-P$). ∎
The result $K^{\prime}(0)=-P$ implies that the infinitesimal change of the modular Hamiltonian with respect to $s$ is exactly generated by the spacetime translation operator $P$. As we vary $s$, the evolution of the operator $K(s)$ is driven by a real geometric translation in spacetime. This confirms that our continuous interpolation path is aligned with an actual spacetime flow: at $s=0$, the direction in which the modular Hamiltonian initially changes is precisely along the generator $P$ (which translates information outward in the emergent space). Thus, the abstract path $\tilde{E}_{s}$ we construct is not just algebraically idempotent, but its tangent direction is that of a physical translation, substantiating the interpretation of the channel as a true spacetime motion of information.
Equipped with $K^{\prime}(0)=-P$ in hand (Lemma 4.1), we can now compute $\tilde{G}$ explicitly and prove the main result.
@@
According to the first-order modular perturbation theory of Araki and Kosaki, the variation of the modular conjugation $J$ under a perturbation of the modular Hamiltonian $K$ satisfies the fundamental identity:
@@
Physically, this factor of 2 arises because $J$ implements a reflection in the modular parameter space; shifting the reflection axis by a unit amount results in a displacement of twice that magnitude for the reflected states.
Substituting the result of Lemma 4.1 ($K^{\prime}(0)=-P$) into this formula, we obtain:
$\tilde{G}\;=\;i\,J_{\tilde{\mathcal{M}}}\,J^{\prime}_{\tilde{\mathcal{N}}}(0)\;=\;-\,2\,K^{\prime}(0)\;=\;-\,2(-P)\;=\;2P~,$ (29)
establishing the algebraic identity on the dense domain $\mathcal{D}_{\text{core}}$. Crucially, this cancellation holds without assuming any special commutation between $K_{\tilde{\mathcal{M}}}$ and $P$: even if $K_{\tilde{\mathcal{M}}},P]\neq 0$, the full perturbative expansion of $J^{\prime}_{\tilde{\mathcal{N}}}(0)$, which typically involves integrals of modular automorphisms, exactly reconstructs the operator $P$ in this canonical setting. Thus, the result $\tilde{G}=2P$ is not an artifact of symmetry, but a robust consequence of the modular perturbation framework.
It remains to lift the equality $\tilde{G}=2P$ from the core domain to the level of closed operators. By construction, the common analytic core $\mathcal{D}_{\text{core}}$ is invariant under the modular flows of both $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$, and hence serves as a joint core for $K_{\tilde{\mathcal{M}}}$, $K_{\tilde{\mathcal{N}}}$, and their difference $P$. On this domain, $2P$ is symmetric (and indeed essentially self-adjoint, since $\mathcal{D}_{\text{core}}$ consists of analytic vectors for $P$). Meanwhile, $\tilde{G}$ is symmetric by definition and is essentially self-adjoint on $\mathcal{D}_{\text{core}}$ by Nelson’s analytic vector theorem (see Lemma 3.10). Therefore, both operators $\tilde{G}$ and $2P$ admit unique self-adjoint extensions which coincide on $\mathcal{D}_{\text{core}}$. By the uniqueness of the self-adjoint closure, we conclude that $\tilde{G}$ and $2P$ are identical as closed operators. This completes the rigorous proof of the identity. ∎
@@ 4.3 Stability Quantification via Non-Commutative $L^{p}$ Spaces
A crucial physical question is whether the identity $\tilde{G}=2P$ is merely an artifact of our idealized mathematical construction at $s=0$, or if it is structurally robust. Specifically, we must ensure that the generator does not fluctuate wildly under small perturbations of the interpolation parameter $s$. We provide a definitive answer using Kosaki’s non-commutative $L^{p}$ theory Kosaki84 ; Kato , which offers the canonical geometric framework for analyzing perturbations in Type II von Neumann algebras.
@@
The generator $\tilde{G}$ is structurally stable. Let $\tilde{G}(s)$ denote the instantaneous generator of the unitary path at parameter $s$. The deviation of its resolvent from the target operator $2P$’s resolvent is quantitatively bounded by the change in the associated projection $e_{\tilde{\mathcal{N}}(s)}$ onto the subalgebra $\tilde{\mathcal{N}}(s)$:
@@
where $||\cdot||$ is the operator norm, and $R_{T}(z)=(T-z)^{-1}$ denotes the resolvent operator.
@@
This confirms that the generator $\tilde{G}(s)$ converges to $2P$ in the generalized strong resolvent sense as $s\to 0$, affirming the structural robustness of the $\tilde{G}=2P$ identity against small deformations. ∎
@@ 4.4 Correlation-Function Test of the Conjecture
The operator identity $\tilde{G}=2P$ suggests a direct connection between the algebraic shift and geometric translation, which can be explored in physical models using correlation functions. Since direct verification of operator-level identities can be challenging in complex quantum gravity models, we propose an observable test using correlation functions.
@@
We conduct this test in the context of the AdS/CFT correspondence. The local algebras $\mathcal{M}$ and $\mathcal{N}$ correspond to subregions of the boundary CFT. The thermofield double state, $|\text{TFD}\rangle$, serves as the canonical purified state holographically dual to the eternal two-sided black hole, representing the entanglement between the Left and Right boundaries. Consequently, let $O_{L}$ and $O_{R}$ be local operators acting on the Left and Right boundaries of the CFT, respectively.
Modeling the teleportation process as a continuous path parameterized by $s$, we define the correlation function $F(s)$ to measure the information transfer:
$F(s):=\langle\text{TFD}|O_{R}\,\Gamma_{s}(O_{L})\,|\text{TFD}\rangle,$ (31)
where the quantum channel is given by $\Gamma_{s}(O_{L})=\tilde{U}(s)O_{L}\tilde{U}(s)^{*}$. Based on our result $\tilde{U}(s)\approx e^{-is\tilde{G}}$, the derivative at $s=0$ (representing the initial teleportation velocity) is predicted to be:
$F^{\prime}(0)=\langle\text{TFD}|O_{R}(-i[\tilde{G},O_{L}])|\text{TFD}\rangle=-2i\langle\text{TFD}|O_{R}[P,O_{L}]|\text{TFD}\rangle.$ (32)
The right-hand side, $2i\langle O_{R}[P,O_{L}]\rangle$, represents the initial velocity of a spacetime translation, scaled by a factor of 2. As established in Section 4.1, this factor of 2 is physically mandatory; it stems from the structure of the canonical shift operator $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$, which is composed of two consecutive modular conjugations. In the TFD geometry, this process effectively doubles the net modular time evolution required to bridge the entanglement wedge of $\mathcal{M}$ to that of $\mathcal{N}$, reflecting the inherent geometric doubling necessary for complete information transfer.
The equality of the two expressions suggests that the algebraic dynamics of teleportation and the geometric dynamics of spacetime translation are two manifestations of the same underlying phenomenon. In the bulk geometry of semiclassical gravity, the generator $P=K_{\mathcal{M}}-K_{\mathcal{N}}$ acts as the operator conjugate to the time shift across the horizon (specifically, the relative modular Hamiltonian). Perturbations generated by $P$ correspond to injecting energy near the horizon, creating a gravitational shock wave that modifies the background geometry. This shock wave induces a physical time delay (or Shapiro delay) for any incoming matter, which is precisely the mechanism governing the correlation function Shenker2014 ; Gao2017 .
In solvable holographic models such as JT gravity or the SYK model, this mechanism manifests explicitly: the algebraic quantity $2i\langle O_{R}[P,O_{L}]\rangle$ corresponds directly to the average time shift experienced by a probe operator $O_{L}$ as it propagates from the left boundary into the bulk, interacts with the shock wave induced by $P$, and is measured by $O_{R}$. For instance, related calculations in these models suggest that $F^{\prime}(0)$ scales proportionally to $\Delta/\beta$, where $\Delta$ is the conformal dimension of $O_{L}$ and $\beta$ is the inverse temperature. This provides a direct, quantitative target for future computational verification, consistent with the framework of information flow and quantum error correction in AdS/CFT Almheiri_2015 ; Penington19 .
Ultimately, this test addresses the core of the black hole information paradox by demonstrating that the protocol for information recovery is physically indistinguishable from a smooth geometric passage through spacetime. Thus, verifying this correlator equality in a concrete holographic setting would lend strong support to the hypothesis that the continuous algebraic shift $\tilde{U}(s)$ is, at least infinitesimally, equivalent to the geometric spacetime translation.

### v3

@@ 4.1 Motivation from Half-Sided Modular Inclusions
Our canonical interpolation constructs a strongly continuous path $\tilde{U}(s)$ that interpolates from the identity to this discrete shift, i.e., $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\tilde{\Gamma}}$. Assuming that this path is generated by a scale-independent operator $\tilde{G}$ (such that $\tilde{U}(s)=e^{-is\tilde{G}}$), the endpoint condition yields:
$e^{-i\tilde{G}}=\tilde{U}(1)=U_{\tilde{\Gamma}}=e^{-2iP}.$ (21)
This exact result in the HSMI limit strongly motivates the identification $\tilde{G}=2P$ for the general case. The factor of 2 in this relationship is dictated by the underlying modular structure. Physically, the canonical shift $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is composed of two consecutive modular conjugations. Since each modular conjugation acts as a geometric reflection (analogous to a CPT inversion), combining two such reflections results in a net translation that is exactly double the elementary geometric shift. This geometric logic justifies the factor of 2 in our operator identity.
4.2 Geometric Derivation via Canonical Path
Our objective is to establish the identity $\tilde{G}=2P$ in general Type III settings, moving beyond the specific symmetry of the HSMI case. Since both $\tilde{G}$ and $2P$ are unbounded self-adjoint operators, a precise proof requires demonstrating that they coincide as closed operators on a common core. We achieve this by employing the modular perturbation theory of Araki, Connes, and Kosaki (see, e.g., Araki:1976zv ; Connes80 ; Kosaki84 ; Kosaki84UC ) and invoking Nelson’s analytic vector theorem Nelson59 to ensure essential self-adjointness.
@@ 4.2 Geometric Derivation via Canonical Path
For the canonical interpolation path $\tilde{\mathcal{N}}(s)$ defined via the Connes–Takesaki cocycle derivative $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$, the modular Hamiltonian $K(s)=K_{\tilde{\mathcal{N}}(s)}$ satisfies the following first-order perturbation relation at $s=0$:
@@
Let $\tilde{\phi}_{0}$ be a faithful normal weight on $\tilde{\mathcal{M}}$. We define the weight $\tilde{\phi}_{1}:=\tilde{\phi}_{0}\circ\tilde{E}$, thereby extending the modular dynamics of the subalgebra $\tilde{\mathcal{N}}$ to the full algebra $\tilde{\mathcal{M}}$. By identifying the modular Hamiltonians via the relation $\Delta=e^{-K}$ (specifically, $\Delta_{\tilde{\mathcal{M}}}=e^{-K_{\tilde{\mathcal{M}}}}$ and $\Delta_{\tilde{\mathcal{N}}}=e^{-K_{\tilde{\mathcal{N}}}}$), the Connes–Takesaki Radon–Nikodym cocycle is explicitly given by:
$D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}\;:=\;\Delta_{\tilde{\phi}_{1}}^{it}\Delta_{\tilde{\phi}_{0}}^{-it}~.$ (23)
The canonical interpolation path $\tilde{\phi}_{s}$ is then constructed via the analytic scaling of this cocycle:
$D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}\;=\;[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{\,st}~.$ (24)
Using the cocycle identity $D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}=\Delta_{s}^{it}\Delta_{0}^{-it}$ (where $\Delta_{0}=\Delta_{\tilde{\mathcal{M}}}$), we can express the modular operator $\Delta_{s}$ along the path as:
$\Delta_{s}^{\,it}\;=\;\big[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}\big]_{st}\,\Delta_{0}^{\,it}\,,\qquad\forall\,t\in\mathbb{R}~.$ (25)
We now compute the derivative of (25) with respect to $s$ at the origin. First, observe that the cocycle $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{u}$ forms a unitary group in the parameter $u=st$, generated by the relative modular Hamiltonian $K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}}$. Applying the chain rule, the derivative of the cocycle term is:
@@
Consequently, differentiating the full expression in (25) yields:
$\frac{d}{ds}\Delta_{s}^{\,it}\Big|_{s=0}\;=\;\left(\frac{d}{ds}\big[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}\big]_{st}\Big|_{s=0}\right)\Delta_{0}^{\,it}\;=\;-it(K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}})\Delta_{0}^{\,it}~.$ (27)
Alternatively, we can evaluate this derivative using the definition $\Delta_{s}^{it}=e^{-itK(s)}$. Applying Duhamel’s formula for the derivative of an exponential operator Kato , a cornerstone of Araki’s modular perturbation theory Araki:1976zv , we obtain:
$\frac{d}{ds}e^{-itK(s)}\Big|_{s=0}\;=\;-i\int_{0}^{t}e^{-i(t-\tau)K_{0}}K^{\prime}(0)e^{-i\tau K_{0}}d\tau~.$ (28)
@@
$K^{\prime}(0)\;=\;K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}}~.$ (29)
Recalling the definition $P:=K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}$, we conclude that $K^{\prime}(0)=-P$. ∎
The result $K^{\prime}(0)=-P$ implies that the infinitesimal change of the modular Hamiltonian with respect to $s$ is governed by the spacetime translation operator $P$. As we vary $s$, the evolution of $K(s)$ is effectively driven by a geometric translation in the emergent spacetime. This confirms that our continuous interpolation path aligns with a physical flow: at $s=0$, the infinitesimal change of the modular Hamiltonian is precisely given by P-P (implying a translation directed opposite to the canonical outward shift). Thus, the abstract path $\tilde{E}_{s}$ is not merely algebraically idempotent; its tangent direction corresponds to a physical translation, substantiating the interpretation of the channel as a geometric transport of information.
Equipped with the result $K^{\prime}(0)=-P$ from Lemma 4.1, we can now explicitly compute $\tilde{G}$ and establish the main theorem.
@@
For a generic perturbation, the variation of $J$ usually involves complex spectral integrals of the modular operator arising from non-commutativity. However, a remarkable feature of the canonical $L^{p}$-interpolation path is that it follows the geodesic of the underlying modular geometry. Consequently, the complex non-commutative terms cancel out, [^2] Explicitly, the path is generated by the analytic scaling $\Delta_{s}^{it}=h^{ist}\Delta_{0}^{it}$. The variation with respect to $s$ extracts the generator $\ln h$ linearly, avoiding the parameter-integral convolution typical of generic perturbations. leaving a simple geometric proportionality:
@@
Physically, this factor of 2 is a direct consequence of the geometric role of $J$. Since the modular conjugation implements a geometric reflection (effectively inverting the emergent spacetime), shifting the reflection axis along the geodesic results in a net displacement of twice that magnitude. This geometric logic is inherent to the canonical $L^{p}$ interpolation and justifies the specific coefficient in our operator identity.
Substituting the result of Lemma 4.1 ($K^{\prime}(0)=-P$) into (31), we obtain:
$\tilde{G}\;=\;-\,2\,K^{\prime}(0)\;=\;-\,2(-P)\;=\;2P~,$ (32)
establishing the algebraic identity on the dense domain $\mathcal{D}_{\text{core}}$. Crucially, this exact reconstruction of $2P$ holds regardless of the commutation relations between $K_{\tilde{\mathcal{M}}}$ and $P$, confirming that the result is a robust consequence of the modular structure rather than an artifact of specific symmetries.
It remains to lift the equality $\tilde{G}=2P$ from the core domain to the level of closed operators. By construction, the common analytic core $\mathcal{D}_{\text{core}}$ is invariant under the modular automorphism groups of both $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$. Consequently, $\mathcal{D}_{\text{core}}$ consists of analytic vectors for the modular Hamiltonians $K_{\tilde{\mathcal{M}}}$ and $K_{\tilde{\mathcal{N}}}$, and thereby for their difference $P$. Thus, $2P$ is essentially self-adjoint on this domain. Similarly, $\tilde{G}$ is essentially self-adjoint on $\mathcal{D}_{\text{core}}$ by Nelson’s analytic vector theorem (Lemma 3.10). Since $\tilde{G}$ and $2P$ coincide on a common core where both are essentially self-adjoint, their unique self-adjoint closures are identical. This completes the proof. ∎
@@ 4.3 Stability Quantification via Non-Commutative $L^{p}$ Spaces
A central physical question is whether the identity $\tilde{G}=2P$ represents a singular coincidence valid only at the boundary $s=0$, or a structurally robust property of the interpolation. Specifically, we must ensure that the generator does not exhibit large or uncontrolled fluctuations under small perturbations of the interpolation parameter $s$. We address this by employing Kosaki’s non-commutative $L^{p}$ theory Kosaki84 ; Kato , which provides the canonical geometric framework for analyzing perturbations in Type II von Neumann algebras.
@@
The generator $\tilde{G}$ exhibits structural stability in the strong resolvent sense. Let $\tilde{G}(s)$ denote the instantaneous generator of the unitary path at parameter $s$. For any fixed complex number $z$ in the resolvent set $\rho(2P)$ (i.e., $z\notin\sigma(2P)$) and any state vector $\xi$ within the dense core domain $\mathcal{D}_{\text{core}}$, the deviation of the resolvent is quantitatively bounded by the variation of the associated projection $e_{\tilde{\mathcal{N}}(s)}$:
@@
Here, $R_{T}(z)=(T-z)^{-1}$ denotes the resolvent operator, and $||\cdot||$ represents the vector norm on the Hilbert space (or the operator norm for the projection difference). The coefficient $C_{z}$ is a stability factor that scales inversely with the distance to the spectrum $\sigma(2P)$ (i.e., $C_{z}\sim\operatorname{dist}(z,\sigma(2P))^{-1}$).
@@
This confirms that the generator $\tilde{G}(s)$ converges to $2P$ in the strong resolvent sense as $s\to 0$, affirming the structural robustness of the $\tilde{G}=2P$ identity against small deformations. ∎
@@ 4.4 Correlation-Function Test of the Conjecture
The operator identity $\tilde{G}=2P$ posits a direct equivalence between the algebraic shift derived from modular theory and the geometric translation of the emergent spacetime. While operator-level identities are mathematically rigid, their physical validity in complex quantum gravity models requires verification through observables. We propose a concrete test using two-point correlation functions within the AdS/CFT correspondence.
@@
We frame this test in the context of a two-sided eternal black hole, holographically dual to the thermofield double state $|\text{TFD}\rangle$. The local algebras $\mathcal{M}$ and $\mathcal{N}$ correspond to nested subregions of the boundary CFT, specifically associated with the Right exterior region. Let $O_{L}$ and $O_{R}$ be local probe operators acting on the Left and Right boundaries, respectively. Since the canonical shift $\tilde{U}(s)$ is generated by the algebra on the Right, it acts non-trivially on $O_{R}$.
To probe whether this algebraic action corresponds to a geometric translation, we define a correlation function $F(s)$ that measures how the Left-Right entanglement implies a correlation change under the algebraic shift of the Right operator:
$F(s)\;:=\;\langle\text{TFD}|\,O_{L}\,\left(\tilde{U}(s)O_{R}\tilde{U}(s)^{*}\right)\,|\text{TFD}\rangle.$ (34)
This correlator probes the “distance” through the wormhole between the fixed Left operator and the algebraically shifted Right operator. Since $\tilde{G}$ is defined as the infinitesimal generator of the path at $s=0$ (satisfying the expansion $\tilde{U}(s)=\mathbf{1}-is\tilde{G}+O(s^{2})$), the initial derivative determines the prediction for the operator identity. If $\tilde{G}=2P$ holds, we obtain:
$F^{\prime}(0)\;=\;\langle\text{TFD}|\,O_{L}\,(-i[\tilde{G},O_{R}])\,|\text{TFD}\rangle\;=\;-2i\langle\text{TFD}|\,O_{L}\,[P,O_{R}]\,|\text{TFD}\rangle.$ (35)
Eq. (35) provides the precise physical interpretation that the algebraic generator $\tilde{G}$ acting on the boundary operator induces a shift exactly twice that induced by the geometric translation generator $P$.
In the bulk gravity description, $P$ generates a geometric time translation along the horizon. Therefore, verifying Eq. (35) amounts to checking if the algebraic shift $\tilde{U}(s)$ effectively displaces the position of the horizon (or the stretched horizon) relative to the probe $O_{L}$ by a coordinate distance of $2s$.
This geometric action is deeply connected to the mechanics of holographic teleportation. In models like the Gao-Jafferis-Wall (GJW) protocol Gao2017 , the traversability is achieved by a gravitational shockwave that shifts the horizon. Here, our result implies that the algebraic generator $\tilde{G}$ explicitly implements this horizon shift without needing an external matter source, effectively simulating the backreaction required to bridge the entanglement wedge.
In semiclassical models like Jackiw-Teitelboim (JT) gravity Saad:2019lba , the quantity $2i\langle O_{L}[P,O_{R}]\rangle$ precisely quantifies this shift, manifesting as a Shapiro time delay for signals crossing the wormhole Shenker2014 ; Maldacena2016 . Thus, the factor of 2 confirms that $\tilde{U}(s)$ physically manipulates the geometry of the stretched horizon, facilitating the information transfer.